A family goes to a school play. Two adult tickets and 1 student ticket costs $8.75. Another family needs 1 adult ticket and 4 student tickets. Their total cost was $10.50. Find the price of each ticket.

A. Write a system of equations for this situation.

B. Solve your system using substitution or elimination. Make sure to show all your work for full credit.

Answer by stanbon(75887)

A family goes to a school play. Two adult tickets and 1 student ticket costs $8.75. Another family needs 1 adult ticket and 4 student tickets. Their total cost was $10.50. Find the price of each
-----
Equations:
2a + s = 875 cents
a + 4s = 1050 cents
-------
Modify for elimination:
2a + s = 875
2a + 8s = 2100
-------
Subtract and solve for "s"::
7s = 1225
s = 175 cents = $1.75 (price of a student ticket)
----------------
Solve for "a"::
2a + s = 875
2a +175 = 875
2a = 700
a = 350 cents = $3.00 (price of an adult ticket)
------------
Cheers,
Stan H.
------------